OpenStudy (anonymous):

solve for the variable : 2/3x = 26

OpenStudy (anonymous):

Multiply 3x on each side.

OpenStudy (anonymous):

divide 26 by 2/3. (multiply by the reciprocal)

OpenStudy (anonymous):

multiplie bouth sides by 3/2: x=26*3/2=39